what mass of Fe(OH) would be produced if 250 ML of 0.25 0M NaOH were added to solution containing excess FeCl3


1
Expert's answer
2021-04-13T08:03:50-0400

Fe(OH)2 + 2NaOH → Na2[Fe(OH)4

Ratio 1 : 2

Moles of NaOH = Volume × Molar Mass

0.25×39.997= 10moles

Moles of Fe(OH) = 10÷2 = 5 moles

Mass =5× 89.86= 434grams

Mass= 434g